Solve for x, and show work....
squareroot of 2 times squareroot of 3 times x = 18

Answer by checkley71(8403)

SQRT2*SQRT3*X=18
XSQRT6=18 SQUARING BOTH SIDES WE GET
6X^2=324
6X^2=324
X^2=54
X=SQRT54
X=7.3484692
########################################
MERRY CHRISTMAS